A man when he travels at 3 kmph he reaches the station  2 mins late. However when he travels at 4 kmph he reaches 2 mins quick. Find  the distance the man has to travel?

Grade:

1 Answers

Anuj K.V. Mishra
10 years ago

   Let his original time of reach be t mins after its start. And req. distance be x meters.

   3 kmph=50 m/min

   4kmph=200/3 m/min

  A/Q

    x = 50(t+2)        -(1)

    x=(200/3)(t-2)    -(2)

    From (1) and (2),we get

    x= 800 m

 

   Hope it helps.